A random sample of size n=613 from a population whose parameter is p=0.43.

  1. What is the mean of the distribution of sample means? Round the answer accurate to 2 decimal places. to find answer

  2. What is the standard deviation of the distribution of sample means? Round the answer accurate to 2 decimal places. to find answer
Answer #1

Solution

Given that,

p = 0.43

1 - p = 1-0.43=0.57

n = 613

mean = = p =0.43

standard deviation = =  [p ( 1 - p ) / n] =   [(0.43*0.57) / 613 ] = 0.02